a 204 inch board is cut into two pieces. one piece is three times the length of the other. find the lengths of the two pieces
Answer by checkley77(12844)

x+3x=204
4x=204
x=204/4
x=51 inches for the shorter piece.
204-51=153 inches for the longer piece.
51+3*51=204
51+153=204
204=204